Let f: [0, 1]→R be a function defined by f(x) = x^m (1-x)^n ,where . m, n∈N


Find the values of m and n such that the Rolle’s Theorem holds for the function


f .

1
Expert's answer
2022-03-20T14:53:31-0400

The function is differentiable on "\left[ 0,1 \right]" , since it is a polynomial.

"f\left( 0 \right) =0,f\left( 1 \right) =0" .

Hence the Rolle’s theorem holds for any "m,n\in \mathbb{N}"